OpenStudy (anonymous):

Factor z 3 - 2z 2 + 9z - 18.

OpenStudy (chealyn98):

z3-2z2+9z-18 Final result : (z2 + 9) • (z - 2) Reformatting the input : Changes made to your input should not affect the solution: (1): "z2" was replaced by "z^2". 1 more similar replacement(s). Step by step solution : Step 1 : Raise z to the 2nd power Exponentiation : Equation at the end of step 1 : (((z3) - (2 • z2)) + 9z) - 18 Step 2 : Raise z to the 3rd power Exponentiation : Equation at the end of step 2 : ((z3 - 2z2) + 9z) - 18 Step 3 : Simplify z3-2z2+9z - 18 Checking for a perfect cube : 3.1 z3-2z2+9z-18 is not a perfect cube Trying to factor by pulling out : 3.2 Factoring: z3-2z2+9z-18 Thoughtfully split the expression at hand into groups, each group having two terms : Group 1: 9z-18 Group 2: z3-2z2 Pull out from each group separately : Group 1: (z-2) • (9) Group 2: (z-2) • (z2) ------------------- Add up the two groups : (z-2) • (z2+9) Which is the desired factorization Polynomial Roots Calculator : 3.3 Find roots (zeroes) of : F(z) = z2+9 Polynomial Roots Calculator is a set of methods aimed at finding values of z for which F(z)=0 Rational Roots Test is one of the above mentioned tools. It would only find Rational Roots that is numbers z which can be expressed as the quotient of two integers The Rational Root Theorem states that if a polynomial zeroes for a rational number P/Q then P is a factor of the Trailing Constant and Q is a factor of the Leading Coefficient In this case, the Leading Coefficient is 1 and the Trailing Constant is 9. The factor(s) are: of the Leading Coefficient : 1 of the Trailing Constant : 1 ,3 ,9 Let us test .... P Q P/Q F(P/Q) Divisor -1 1 -1.00 10.00 -3 1 -3.00 18.00 -9 1 -9.00 90.00 1 1 1.00 10.00 3 1 3.00 18.00 9 1 9.00 90.00 Polynomial Roots Calculator found no rational roots Final result : (z2 + 9) • (z - 2)

OpenStudy (jango_in_dtown):

z^2(z-2)+9(z-2)=(z^2+9)(z-2)=(z+3i)(z-3i)(z-2)

OpenStudy (chealyn98):

http://www.tiger-algebra.com/drill/z3-2z2_9z-18/